OpenStudy (anonymous):

plz explain this statement: simple harmonic motion is the projection of uniform circular motion onto a diameter?

OpenStudy (anonymous):

 suppose a particle is moving on a circular path of radius A with constant angular speed w. assigning theta =0 at x=0. then theta at time t = w*t now projection of position vector of particle on X-axix at time t= A cos(theta) = A cos (w*t)

OpenStudy (anonymous):

@celos why t=w*t?

OpenStudy (anonymous):

sorry, i mean at time t, theta = w*t

OpenStudy (anonymous):

yes but why we take theta = w*t?

OpenStudy (anonymous):

angular displacement= angular velocity*time let at time T=0 position of particle is at (A,0) so, at time T=t theta=w*t

OpenStudy (anonymous):

@celos thanx